# Distribution-free data-driven smooth tests without $\chi^{2}$

### v2

@@ 3.1 Empirical illustrations of smooth tests via projection
In what follows, the parameters $\theta$ and $\sigma$ are set equal to -10 and 2, respectively, and are assumed to be known.
A dataset of $n=100$ observations is generated from the density specified in (16), with the true value of the asymmetry parameter set to be $\bm{\beta}_{0}=0.1$. The maximum likelihood estimator $\widehat{\bm{\beta}}_{n}$ of $\bm{\beta}$ based on these observations is obtained by solving the following equation (Kotz et al.,, 2002):
@@
Given the complexity of solving (17), it is anticipated that using the projected bootstrap – in which the estimator is obtained only once rather than recalculated on each bootstrap sample – will expedite the simulation procedure.
To test the hypothesis in (2), we consider the order selection statistic in (7) and the subset selection statistic in (9). Their null distributions are simulated via three methods: the classical parametric bootstrap, the projected bootstrap, and the Monte Carlo method with data generated from the true distribution. By incorporating a simulation of Monte Carlo samples, we aim to determine whether the two bootstrap procedures considered can accurately recover the true null distribution of the test statistics for a sample size of $n=100$. All simulations are conducted using $100,000$ replicates. The orthonormal basis functions $\{h_{j\bm{\beta}}\}_{j=1}^{M_{n}}$ are chosen as compositions of the normalized shifted Legendre polynomials $h_{j}$ on $[0,1]$ with the CDF $G_{\bm{\beta}}$, i.e., $h_{j\bm{\beta}}=h_{j}\circ G_{\bm{\beta}}$. The maximum number of basis functions, $M_{n}$, is set to be 10.
Figure 1: Comparing the simulated null distributions of the order selection test statistic in (7) (Left), and subset selection test statistic in (9) (Right) via the parametric bootstrap (orange dotted lines), the projected bootstrap (grey dashed lines), and the Monte Carlo method (blue solid lines).
The results of the simulation, presented in Figure 1, demonstrate that the simulated null distributions of the order selection and subset selection test statistics are consistent across all three methods considered. This indicates that both bootstrap procedures successfully recover the true null distribution of the test statistics, even with a sample size as small as 100 observations. In terms of computational efficiency, as expected, the projected bootstrap method is significantly faster than the classical parametric bootstrap. Specifically, simulating the null distributions of the order selection and subset selection statistics using 100,000 replicates required approximately 12 minutes of CPU time using the projected bootstrap, less than a fourth of the time needed by the classical parametric bootstrap, which required approximately 51 minutes.
@@ 4 Distribution-free smooth tests via K2 transform
In general, statistics given by functionals of the process $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ are not distribution-free. Thus, while the projected bootstrap introduced in Section 3 can reduce the computational burden of simulating their null distribution, a different simulation must be implemented for each model being tested.
@@
This section demonstrates that another route for distribution-freeness is available to practitioners. Instead of combining the values of the process $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ in a rather specific – and possibly forceful – manner to construct at most a few distribution-free statistics, the distribution-free property can be retrieved by ensuring that the empirical process $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ is itself asymptotically distribution-free under the null. Such an approach guarantees that the limiting null distribution of all its functionals is also distribution-free, thereby providing the user with an entire family of distribution-free statistics. The tool that enables such a construction is the so-called Khmaladze-2 (K2) transform [^2] The name “Khmaladze-2 transform” is used to distinguish the transformation employed in this manuscript and proposed in Khmaladze, (2016) from an earlier “Khmaladze transform” introduced by the same author in Khmaladze, (1982). introduced by Khmaladze, (2016).
@@
In classical nonparametric goodness-of-fit testing, the probability integral transform is the transformation commonly employed to map the empirical process into the uniform empirical process, which is known to be asymptotically distributed as a standard Brownian bridge. Likewise, in the parametric setting, the K2 transform allows us to map a wide range of different projected empirical processes into standard projections that share the same limiting null distribution. When applied in the context of smooth tests, the K2 transform implicitly guides the construction of an orthonormal basis for $L^{2}(G_{\bm{\beta}})$ that guarantees the distribution-freeness of the process indexed by functions in such a basis and all its functionals.
@@
Let $F_{\bm{\gamma}}$ be a “reference distribution” defined on the same support $\mathcal{X}$ as $G_{\bm{\beta}}$. For the moment, assume that $\bm{\gamma}$ and $\bm{\beta}$ share the same dimension; an extension to the case where $\bm{\gamma}$ has fewer parameters than $\bm{\beta}$ will be discussed in Section 4.1. The distribution $F_{\bm{\gamma}}$ constitutes the starting point of our construction and should be chosen to be simple – that is, easy to simulate from, differentiate, and evaluate. Consider the Hilbert space $L^{2}(F_{\bm{\gamma}})=\{\phi_{\bm{\gamma}}:\left\langle\phi_{\bm{\gamma}},\phi_{\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}<\infty\}$, with inner product
$\left\langle\phi_{\bm{\gamma}},\phi^{\prime}_{\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}=\int_{\mathcal{X}}\phi_{\bm{\gamma}}(x)\phi^{\prime}_{\bm{\gamma}}(x)dF_{\bm{\gamma}}(x),$
@@
and let $\bm{a}_{\bm{\gamma}}=[a_{\gamma_{1}},\cdots,a_{{\gamma_{p}}}]^{T}$ be the orthonormalized score function of $F_{\bm{\gamma}}$. Denote with $\mathcal{L}\left(F_{\bm{\gamma}}\right)$ and $\mathcal{L}_{\perp}\left(F_{\bm{\gamma}}\right)$ the subspaces of $L^{2}\left(F_{\bm{\gamma}}\right)$ given by
@@
$\mathcal{L}\left(F_{\bm{\gamma}}\right)=\left\{\phi_{\bm{\gamma}}\in L^{2}\left(F_{\bm{\gamma}}\right):\left\langle\phi_{\bm{\gamma}},\bm{1}\right\rangle_{F_{\bm{\gamma}}}=0\right\},\ \ \ \mathcal{L}_{\perp}\left(F_{\bm{\gamma}}\right)=\left\{\widetilde{\phi}_{\bm{\gamma}}\in\mathcal{L}\left(F_{\bm{\gamma}}\right):\bigl{\langle}{\widetilde{\phi}_{\bm{\gamma}},\bm{a}_{\bm{\gamma}}\bigl{\rangle}}_{F_{\bm{\gamma}}}=0\right\}.$
@@
For any given orthonormal basis $\{\phi_{j\bm{\gamma}}\}_{j=1}^{\infty}$ in $\mathcal{L}(F_{\bm{\gamma}})$ the empirical process indexed by the residuals
$\widetilde{\phi}_{j\bm{\gamma}}=\phi_{j\bm{\gamma}}-\bm{a}^{T}_{\bm{\gamma}}\left\langle\bm{a}^{T}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}},$
@@
respectively. Such a limiting distribution is the “standard” distribution that will be recovered through the K2 transform. In particular, since a Gaussian process is fully characterized by its mean and covariance, the K2 transform enables the construction of an orthonormal basis $\{h_{j\bm{\beta}}\}_{j=1}^{\infty}$ for $L^{2}(G_{\bm{\beta}})$, hereinafter referred to as K2 orthonormal basis, whose residuals, $\{\widetilde{h}_{j\bm{\beta}}\}_{j=1}^{\infty}$ satisfy
@@
It follows that the processes $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ and $v_{F,n}(\widetilde{\phi}_{j\bm{\gamma}})$ – as well as their functionals – have the same limiting distribution under $G_{\bm{\beta}}$ and $F_{\bm{\gamma}}$, respectively, thereby enabling the construction of a large class of asymptotically distribution-free statistics for testing (2).
For example, let $\widehat{\bm{\beta}}_{n}$ and $\widehat{\bm{\gamma}}_{n}$ denote the estimators of $\bm{\beta}$ and $\bm{\gamma}$, respectively. Consider the order-selection and subset-selection statistics defined in (7) and (9), in which $\widehat{S}_{m,n}$ chosen to be the (unnormalized) generalized score statistic in (6), with $\widetilde{h}_{j\widehat{\bm{\beta}}_{n}}$ corresponding to the residuals of the K2 orthonormal basis. Under $G_{\bm{\beta}}$, such statistics have the same limiting distribution as the statistics
$\max_{{m\in\mathcal{M}_{n}}}\sum_{j=1}^{m}\frac{v^{2}_{F,n}(\widetilde{\phi}_{j\widehat{\bm{\gamma}}_{n}})}{m}\ \ \text{ and }\ \ \max_{{B\subseteq\mathcal{M}_{n}:B\neq\emptyset}}\sum_{j\in B}\frac{v^{2}_{F,n}(\widetilde{\phi}_{j\widehat{\bm{\gamma}}_{n}})}{|B|},$ (18)
@@ 4.1 On the construction of the K2 orthonormal basis
It can be easily verified that for all $j,$
@@
It follows that the functions $l_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$ have the same covariance as $\phi_{j\bm{\gamma}}$ under $F_{\bm{\gamma}}$, but do not share the same mean. Thus, they belong to $L^{2}(G_{\bm{\beta}})$ but not to $\mathcal{L}(G_{\bm{\beta}})$.
@@
The operator $K$ satisfying the above requirements can be the reflection
@@
$Kl_{\bm{\gamma},\bm{\beta}}=\bm{1},\quad K\bm{1}=l_{\bm{\gamma},\bm{\beta}},\quad\text{and if }\phi\perp span(\bm{1},l_{\bm{\gamma},\bm{\beta}}),\quad K\phi=\phi.$
Thus far, we have demonstrated that the functions $Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\in\mathcal{L}(G_{\bm{\beta}})$ have the same mean and covariance as the functions $\phi_{j\bm{\gamma}}\in\mathcal{L}(F_{\bm{\gamma}})$. Therefore, when $\bm{\gamma}$ and $\bm{\beta}$ are known, test statistics can be constructed as functionals of the empirical process $v_{G,n}(h_{j\bm{\beta}})$ and they are asymptotically distribution-free if $h_{j\bm{\beta}}=Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$.
When $\bm{\gamma}$ and $\bm{\beta}$ are unknown, an additional step is necessary to map functions $\widetilde{\phi}_{j\bm{\gamma}}\in\mathcal{L}_{\perp}(F_{\bm{\gamma}})$ into functions $\widetilde{h}_{j\bm{\beta}}$ that share the same mean and covariance. For instance, if we were to naively choose $\widetilde{h}_{j\bm{\beta}}$ to be
$\widetilde{h}_{j\bm{\beta}}=Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}-\bm{b}^{T}_{\bm{\beta}}\left\langle\bm{b}^{T}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}},$
@@
$\langle\widetilde{h}_{j\bm{\beta}},\widetilde{h}_{s\bm{\beta}}\rangle_{G_{\bm{\beta}}}=\left\langle\phi_{j\bm{\gamma}},\phi_{s\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}-\left\langle\bm{b}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}\left\langle\bm{b}^{T}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{s\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}\neq\langle\widetilde{\phi}_{j\bm{\gamma}},\widetilde{\phi}_{s\bm{\gamma}}\rangle_{F_{\bm{\gamma}}}.$
@@
$\big{\langle}{\bm{b}^{T}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\big{\rangle}}_{G_{\bm{\beta}}}=\left\langle\bm{a}^{T}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}.$
This motivates us to seek an operator, $\bm{U}_{p}$, such that, when applied to the functions $Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$, we obtain
$\left\langle\bm{b}^{T}_{\bm{\beta}},\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}=\left\langle\bm{a}^{T}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}},$ (21)
while ensuring their means and covariances are unchanged, that is,
@@
Let $c_{\bm{\lambda}_{k}}$ be the kthkth component of the vector-valued function $Kl_{\bm{\gamma},\bm{\beta}}\bm{a}_{\bm{\gamma}}$, i.e.,
@@
While any unitary operator $\bm{U}_{p}$ such that $\bm{U}_{p}\bm{1}=\bm{1}$ fulfills the requirements in (22), to ensure (21) holds, we need to choose $\bm{U}_{p}$ so that $\bm{U}_{p}c_{\bm{\lambda}_{k}}=b_{\bm{\beta}_{k}}$, for each $k=1,\dots,p$.
@@
$U_{b_{\bm{\beta}_{2}}c_{\lambda_{2}}}U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{1}}=U_{b_{\bm{\beta}_{2}}c_{\lambda_{2}}}b_{\bm{\beta}_{1}},$
@@
$U_{b_{\bm{\beta}_{2}}c_{\lambda_{2}}}b_{\bm{\beta}_{1}}\neq b_{\bm{\beta}_{1}}\quad\text{unless}\quad c_{\lambda_{2}}\perp b_{\bm{\beta}_{1}}.$
@@
$\widetilde{c}_{\lambda_{k}}$ $=U_{b_{\bm{\beta}_{k-1}}\widetilde{c}_{\lambda_{k-1}}}\cdots U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{k}},\quad k=2,\ldots,p.$ (23)
@@
$\bm{U}_{p}=U_{b_{\bm{\beta}_{p}}\widetilde{c}_{\lambda_{p}}}\cdots U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}.$
Since $\bm{U}_{p}$ is a composition of unitary operators on $L^{2}\left(G_{\bm{\beta}}\right)$, it is also unitary on $L^{2}\left(G_{\bm{\beta}}\right)$. Moreover,
$\bm{U}_{p}c_{\lambda_{k}}=U_{b_{\bm{\beta}_{p}}\widetilde{c}_{\lambda_{p}}}\cdots U_{b_{\bm{\beta}_{k}}\widetilde{c}_{\lambda_{k}}}\widetilde{c}_{\lambda_{k}}=U_{b_{\bm{\beta}_{p}}\widetilde{c}_{\lambda_{p}}}\cdots U_{b_{\bm{\beta}_{k+1}}\widetilde{c}_{\lambda_{k+1}}}b_{\bm{\beta}_{k}}=b_{\bm{\beta}_{k}},$

### v3

@@ 3.1 Empirical illustrations of smooth tests via the projected bootstrap
In what follows, the parameters $\theta$ and $\sigma$ are set equal to 0 and 2, respectively, and are assumed to be known.
Figure 1: Comparing the simulated null distributions of the order selection test statistic in (7) (Left), and subset selection test statistic in (9) (Right) via the parametric bootstrap (orange dotted lines), the projected bootstrap (grey dashed lines), and Monte Carlo (blue solid lines).
A dataset $\bm{x}_{n}=(x_{1},\dots,x_{n})$ of $n=100$ observations is generated from the density specified in (16), with the true value of the asymmetry parameter set to be $\bm{\beta}_{0}=0.1$. In the present simulation study, such a dataset plays the same role as the observed data sample in a real-data analysis. The maximum likelihood estimator $\widehat{\bm{\beta}}_{n}$ of $\bm{\beta}$ based on $\bm{x}_{n}$ is obtained by solving the following equation (Kotz et al., 2002):
@@
Given the complexity of finding the solution of (17), it is anticipated that using the projected bootstrap – in which the estimator is obtained only once rather than recalculated on each bootstrap sample – will expedite the simulation procedure.
To test the hypothesis in (2), we consider the order selection statistic in (7) and the subset selection statistic in (9). Their null distributions are simulated via three methods: the classical parametric bootstrap, the projected bootstrap, and Monte Carlo with data generated from the true distribution. By incorporating a simulation of Monte Carlo samples, we aim to determine whether the two bootstrap procedures considered can accurately recover the true null distribution of the test statistics for a sample size of $n=100$. All simulations are conducted using $100,000$ replicates. The orthonormal basis functions $\{h_{j\bm{\beta}}\}_{j=1}^{M_{n}}$ are chosen as compositions of the normalized shifted Legendre polynomials $h_{j}$ on $[0,1]$ with the CDF $G_{\bm{\beta}}$, i.e., $h_{j\bm{\beta}}=h_{j}\circ G_{\bm{\beta}}$. The maximum number of basis functions $M_{n}$ considered is 10.
The results of the simulation, presented in Figure 1, demonstrate that the simulated null distributions of the order selection and subset selection test statistics are consistent across all three methods considered. This indicates that both bootstrap procedures successfully recover the true null distribution of the test statistics, even with a sample size as small as 100 observations. In terms of computational efficiency, however, the projected bootstrap method is significantly faster than the classical parametric bootstrap. Specifically, simulating the null distributions of the order selection and subset selection statistics using 100,000 replicates required approximately 9 minutes of CPU time using the projected bootstrap, less than a fourth of the time needed by the classical parametric bootstrap, which required approximately 37 minutes.
The p-values for testing (2) are also computed by comparing the value of the order selection and subset selection statistics calculated on $\bm{x}_{n}$ with their null distribution simulated by means of the projected bootstrap. For the order selection statistic, the p-value obtained is 0.843, and for the subset selection statistic, it amounts to 0.963.
@@ 4 Distribution-free smooth tests via K2 transform
In general, statistics given by functionals of the process $v_{\bm{\beta},n}(\widetilde{h}_{j\bm{\beta}})$ are not distribution-free. Thus, while the projected bootstrap introduced in Section 3 can reduce the computational burden of simulating their null distribution, a different simulation must be implemented for each model being tested.
@@
This section demonstrates that another route to distribution-freeness is available to practitioners. Instead of combining the values of the process $v_{\bm{\beta},n}(\widetilde{h}_{j\bm{\beta}})$ in a rather specific – and possibly forceful – manner to construct, at most, a few distribution-free statistics, the distribution-free property can be retrieved by ensuring that the empirical process $v_{\bm{\beta},n}(\widetilde{h}_{j\bm{\beta}})$ is itself asymptotically distribution-free under the null. Such an approach guarantees that the limiting null distribution of all its functionals is also distribution-free, thereby providing the user with an entire family of asymptotically distribution-free statistics. The tool that enables such a construction is the so-called Khmaladze-2 (K2) transform introduced by Khmaladze (2016).
@@
In classical nonparametric goodness-of-fit testing, the probability integral transform is the transformation commonly employed to map the empirical process into the uniform empirical process, which is known to be asymptotically distributed as a standard Brownian bridge. Distribution-freeness is therefore achieved by means of a simple change of variable. One can think of the K2 transform also as a change of variable, but in functional space. In the parametric setting, such a change of variable allows us to retrieve asymptotic distribution-freeness by mapping a wide range of different projected Brownian motions, arising when estimating different distributions, into the same “standard” projected Brownian motion of choice. Moreover, when applied in the context of smooth tests, the K2 transform implicitly guides the construction of an orthonormal basis for $L^{2}(G_{\bm{\beta}})$ that guarantees the distribution-freeness of the empirical process indexed by functions in such a basis, as well as that of all its functionals.
@@
Let $F_{\bm{\gamma}}$ be a “reference distribution” defined on the same support $\mathcal{X}$ as $G_{\bm{\beta}}$. For the moment, assume that $\bm{\gamma}$ and $\bm{\beta}$ share the same dimension; an extension to the case where $\bm{\gamma}$ has fewer parameters than $\bm{\beta}$ will be discussed in Section 4.1. The distribution $F_{\bm{\gamma}}$ constitutes the starting point of our construction and should be chosen to be simple – that is, easy to simulate from, differentiate, and evaluate. Consider the Hilbert space $L^{2}(F_{\bm{\gamma}})=\{\phi:\left\langle\phi,\phi\right\rangle_{F_{\bm{\gamma}}}<\infty\}$, with inner product
$\left\langle\phi,\phi^{\prime}\right\rangle_{F_{\bm{\gamma}}}=\int_{\mathcal{X}}\phi(x)\phi^{\prime}(x)dF_{\bm{\gamma}}(x),$
and the empirical process
$v_{\gamma,n}(\phi)=\int_{\mathcal{X}}\phi(x)dv_{\bm{\gamma},n}(x)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\bigl[\phi(X_{i})-\mathbb{E}_{F_{\bm{{\gamma}}}}[\phi(X_{i})]\bigl]$
@@
indexed by such functions. Let $\bm{a}_{\bm{\gamma}}=[a_{\gamma_{1}},\cdots,a_{{\gamma_{p}}}]^{T}$ be the orthonormalized score function of $F_{\bm{\gamma}}$. Denote with $\mathcal{L}\left(F_{\bm{\gamma}}\right)$ and $\mathcal{L}_{\perp}\left(F_{\bm{\gamma}}\right)$ the subspaces of $L^{2}\left(F_{\bm{\gamma}}\right)$ given by
@@
$\mathcal{L}\left(F_{\bm{\gamma}}\right)=\left\{\phi\in L^{2}\left(F_{\bm{\gamma}}\right):\left\langle\phi,\bm{1}\right\rangle_{F_{\bm{\gamma}}}=0\right\},\ \ \ \mathcal{L}_{\perp}\left(F_{\bm{\gamma}}\right)=\left\{\widetilde{\phi}\in\mathcal{L}\left(F_{\bm{\gamma}}\right):\bigl\langle{\widetilde{\phi},\bm{a}_{\bm{\gamma}}\bigl\rangle}_{F_{\bm{\gamma}}}=0\right\}.$
@@
For any given orthonormal basis $\{\phi_{j\bm{\gamma}}\}_{j=1}^{\infty}$ of $L^{2}(F_{\bm{\gamma}})$, with elements in $\mathcal{L}(F_{\bm{\gamma}})$, the empirical process $v_{\gamma,n}(\widetilde{\phi}_{j\bm{\gamma}})$ indexed by the residuals
$\widetilde{\phi}_{j\bm{\gamma}}=\phi_{j\bm{\gamma}}-\bm{a}^{T}_{\bm{\gamma}}\left\langle\bm{a}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}},$
@@
respectively. Such a limiting distribution is the “standard” distribution that will be recovered through the K2 transform. In particular, since a Gaussian process is fully characterized by its mean and covariance functions, the K2 transform enables the construction of an orthonormal basis $\{h_{j\bm{\beta}}\}_{j=1}^{\infty}$ for $L^{2}(G_{\bm{\beta}})$, hereinafter referred to as K2 orthonormal basis, whose residuals, $\{\widetilde{h}_{j\bm{\beta}}\}_{j=1}^{\infty}$ satisfy
@@
It follows that the processes $v_{\bm{\beta},n}(\widetilde{h}_{j\bm{\beta}})$ and $v_{\bm{\gamma},n}(\widetilde{\phi}_{j\bm{\gamma}})$, as well as all their functionals, have the same limiting distribution under $G_{\bm{\beta}}$ and $F_{\bm{\gamma}}$, respectively, thereby enabling the construction of a large class of asymptotically distribution-free statistics for testing (2).
For example, let $\widehat{\bm{\beta}}_{n}$ and $\widehat{\bm{\gamma}}_{n}$ denote the estimators of $\bm{\beta}$ and $\bm{\gamma}$, respectively. Consider the order-selection and subset-selection statistics defined in (7) and (9), in which $\widehat{S}_{m,n}$ is chosen to be the (unnormalized) generalized score statistic in (6), with $\widetilde{h}_{j\widehat{\bm{\beta}}_{n}}$ corresponding to the residuals of the K2 orthonormal basis. Under $G_{\bm{\beta}}$, such statistics have the same limiting distribution as the statistics
$\max_{{m\in\mathcal{M}_{n}}}\sum_{j=1}^{m}\frac{v^{2}_{\widehat{\bm{\gamma}}_{n},n}(\widetilde{\phi}_{j\widehat{\bm{\gamma}}_{n}})}{m}\ \ \text{ and }\ \ \max_{{B\subseteq\mathcal{M}_{n}:B\neq\emptyset}}\sum_{j\in B}\frac{v^{2}_{\widehat{\bm{\gamma}}_{n},n}(\widetilde{\phi}_{j\widehat{\bm{\gamma}}_{n}})}{|B|},$ (18)
@@ 4.1 On the construction of the K2 orthonormal basis
It can be easily verified that for all $j$’s,
@@
It follows that the functions $l_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$ have the same covariance as $\phi_{j\bm{\gamma}}$ under $F_{\bm{\gamma}}$, but do not have the same mean. Thus, they belong to $L^{2}(G_{\bm{\beta}})$ but not to its subspace $\mathcal{L}(G_{\bm{\beta}})$.
@@
A unitary operator $K$ satisfying the above requirements is the reflection
@@
$Kl_{\bm{\gamma},\bm{\beta}}=\bm{1}\quad\text{and}\quad K\bm{1}=l_{\bm{\gamma},\bm{\beta}}.$
Moreover, for any $\zeta\in L^{2}(G_{\bm{\beta}})$ such that $\left\langle\zeta,l_{\bm{\gamma},\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=\left\langle\zeta,\bm{1}\right\rangle_{G_{\bm{\beta}}}=0$, $K\zeta=\zeta.$
Thus far, we have demonstrated that the functions $Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\in\mathcal{L}(G_{\bm{\beta}})$ have the same mean and covariance as the functions $\phi_{j\bm{\gamma}}\in\mathcal{L}(F_{\bm{\gamma}})$. Therefore, when $\bm{\beta}$ is known, test statistics constructed as functionals of the empirical process $v_{\bm{\beta},n}(h_{j\bm{\beta}})$ are asymptotically distribution-free if $h_{j\bm{\beta}}=Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$.
When $\bm{\gamma}$ and $\bm{\beta}$ are unknown, an additional step is necessary to map functions $\widetilde{\phi}_{j\bm{\gamma}}\in\mathcal{L}_{\perp}(F_{\bm{\gamma}})$ into functions $\widetilde{h}_{j\bm{\beta}}\in\mathcal{L}_{\perp}(G_{\bm{\beta}})$ with the same mean and covariance. For instance, if we were to naively choose $\widetilde{h}_{j\bm{\beta}}$ to be
$\widetilde{h}_{j\bm{\beta}}=Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}-\bm{b}^{T}_{\bm{\beta}}\left\langle\bm{b}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}},$
@@
$\langle\widetilde{h}_{j\bm{\beta}},\widetilde{h}_{s\bm{\beta}}\rangle_{G_{\bm{\beta}}}=\left\langle\phi_{j\bm{\gamma}},\phi_{s\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}-\left\langle Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},\bm{b}^{T}_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}\left\langle\bm{b}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{s\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}\neq\langle\widetilde{\phi}_{j\bm{\gamma}},\widetilde{\phi}_{s\bm{\gamma}}\rangle_{F_{\bm{\gamma}}}.$
@@
$\big\langle{\bm{b}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\big\rangle}_{G_{\bm{\beta}}}=\left\langle\bm{a}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}$
for all $j$’s. This motivates the construction of an operator, $\bm{U}_{p}$, such that, when applied to the functions $Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$, guarantees
$\left\langle\bm{b}_{\bm{\beta}},\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}=\left\langle\bm{a}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}},$ (21)
and at the same time preserves the mean and covariance structure, i.e.,
@@
Let $c_{\bm{\lambda}_{k}}$ be the $k$th component of the vector function $Kl_{\bm{\gamma},\bm{\beta}}\bm{a}_{\bm{\gamma}}$, i.e.,
@@
While any unitary operator $\bm{U}_{p}$ such that $\bm{U}_{p}\bm{1}=\bm{1}$ fulfills the requirements in (22), to ensure (21) holds, $\bm{U}_{p}$ must be choosen so that $\bm{U}_{p}c_{\bm{\lambda}_{k}}=b_{\bm{\beta}_{k}}$, for each $k=1,\dots,p$.
@@
$U_{b_{\bm{\beta}_{2}}c_{\lambda_{2}}}\circ U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{1}}=U_{b_{\bm{\beta}_{2}}c_{\lambda_{2}}}b_{\bm{\beta}_{1}},$
@@
$U_{b_{\bm{\beta}_{2}}c_{\lambda_{2}}}b_{\bm{\beta}_{1}}\neq b_{\bm{\beta}_{1}}\quad\text{unless}\quad\left\langle c_{\lambda_{2}},b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=0.$
@@
$\widetilde{c}_{\lambda_{k}}$ $=U_{b_{\bm{\beta}_{k-1}}\widetilde{c}_{\lambda_{k-1}}}\circ\cdots\circ U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{k}},\quad k=2,\ldots,p.$ (23)
@@
$\bm{U}_{p}=U_{b_{\bm{\beta}_{p}}\widetilde{c}_{\lambda_{p}}}\circ\cdots\circ U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}.$
Since $\bm{U}_{p}$ is a composition of unitary operators on $L^{2}\left(G_{\bm{\beta}}\right)$, it is itself unitary on $L^{2}\left(G_{\bm{\beta}}\right)$. Moreover,
$\bm{U}_{p}c_{\lambda_{k}}=U_{b_{\bm{\beta}_{p}}\widetilde{c}_{\lambda_{p}}}\circ\cdots\circ U_{b_{\bm{\beta}_{k}}\widetilde{c}_{\lambda_{k}}}\widetilde{c}_{\lambda_{k}}=U_{b_{\bm{\beta}_{p}}\widetilde{c}_{\lambda_{p}}}\circ\cdots\circ U_{b_{\bm{\beta}_{k+1}}\widetilde{c}_{\lambda_{k+1}}}b_{\bm{\beta}_{k}}=b_{\bm{\beta}_{k}},$